On August 31, the balance sheet of Marigold Corp. showed Cash $12,000, Accounts Receivable $4,700, Supplies $600, Equipment $6,000, Accounts Payable $6,600, Common Stock $16,050, and Retained Earnings $650. During September, the following transactions occurred.

1. Paid $3,500 cash for accounts payable due.
2. Collected $2,050 of accounts receivable.
3. Purchased additional equipment for $2,350, paying $900 in cash and the balance on account.
4. Recognized revenue of $7,900, of which $2,550 is collected in cash and the balance is due in October.
5. Declared and paid a $2,250 cash dividend.
6. Paid salaries $2,100, rent for September $1,150, and advertising expense $100.
7. Incurred utilities expense for month on account $180.
8. Received $12,000 from Capital Bank on a 6-month note payable.

a. Prepare a tabular analysis of the September transactions beginning with August 31 balances. (If a transaction causes a decrease in Assets, Liabilities or Stockholders' Equity, place a negative sign (or parentheses) in front of the amount entered for the particular Asset, Liability or Equity item that was reduced. See Illustration 1-8 for example.)


c. Prepare a retained earnings statement for September. (List items that increase retained earnings first.)b. Prepare an income statement for September.


d. Prepare a balance sheet at September 30. (List Assets in order of liquidity.)
Answer #1

a.

Transaction Balance Sheet Income Statement
Assets = Liabilities + Stockholders' Equity Revenue - Expense = Net Income
Cash Accounts Receivable Supplies Equipment Accounts Payable Note Payable Common Stock Retained Earnings
Aug. 31 balance 12000 4700 600 6000 = 6600 + 16050 650 - =
1 -3500 = -3500 + - =
2 2050 -2050 = + - =
3 -900 2350 = 1450 + - =
4 2550 5350 = + 7900 7900 - = 7900
5 -2250 = + -2250 - =
6a. -2100 = + -2100 - 2100 = -2100
6b. -1150 = + -1150 - 1150 = -1150
6c. -100 = + -100 - 100 = -100
7 = 180 + -180 - 180 = -180
8 12000 = 12000 + - =
Sep. 30 balance 18600 8000 600 8350 = 4730 12000 + 16050 2770 7900 - 3530 = 4370

b.

Marigold Corp.
Income Statement
For the Month Ended September
Revenue 7900
Less Expenses:
Salaries expense 2100
Rent expense 1150
Advertising expense 100
Utilities expense 180
Total expense 3530
Net income $ 4370

c.

Marigold Corp.
Statement of Retained Earnings
For the Month Ended September
Balance, August 31 650
Add: Net income 4370
5020
Less: Dividends 2250
Balance, September 30 2770

d.

Marigold Corp.
Balance Sheet
September 30
Assets
Cash 18600
Accounts receivable 8000
Supplies 600
Equipment 8350
Total assets 35550
Liabilities and Stockholders' Equity
Notes payable 12000
Accounts payable 4730
Total liabilities 16730
Common stock 16050
Retained earnings 2770
Total stockholders' equity 18820
Total liabilities and stockholders' equity 35550